OpenStudy (anonymous):

The quotient of twice a number and five is negative two. Find the number. Translate the word problem to an equation. Which steps describe how to solve the equation? Multiply by 5; then, multiply by 1/2. Multiply by 5; then, divide by 1/2. Divide by 5; then, multiply by 1/2. Subtract 5; then, multiply by -2.

OpenStudy (misssunshinexxoxo):

Which do you believe it is?

OpenStudy (muscrat123):

yes. which do u believe it is-its ok if ur wrong

OpenStudy (misssunshinexxoxo):

2x/5=-2 cross multiply. 2x=-10 x=-10/2 x=-5 ans. 2(-5)/5=-2 -10/5=-2 -2=-2

OpenStudy (anonymous):

i think its c

OpenStudy (anonymous):

why?

OpenStudy (misssunshinexxoxo):

2x/5=-2 cross multiply. 2x=-10 x=-10/2 x=-5 ans. 2(-5)/5=-2 -10/5=-2 -2=-2

OpenStudy (misssunshinexxoxo):

Yes you are right my bad typo; C is right

OpenStudy (muscrat123):

lol

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thanks